Interim Constitution of the SHIP Collaboration – Appendix A

(7 October 2015)

Members of the Conveners Meeting, previously named interim Steering Group

O. Boiarskyi W. Bonivento G. De Lellis H. Dijkstra V. Egorychev M. Ferro-Luzzi A. Golutvin D. Gorbunov E. van Herwijnen R. Jacobsson H. Lacker G. Lanfranchi A. Malinin M. Patel F. Rademakers O. Ruchayskiy T. Ruf N. Serra B. Storaci M. Shaposhnikov A. Ustyuzhanin M. Villa